Grade 11Mechanics

how to solve sum
if a,b,c be the distances travelled by the body during xth,yth and zth second from start , then prove that
a(y-z) +b(z-x) +c(x-y) = 0

PRANAV SATHEESH
11 Years agoGrade 11

1 Answer

Adarsh
ApprovedApproved Tutor Answer11 Years ago
According to the question
a= u + A/2(2x-1)
b= u + A/2(2y-1)
c= u + A/2(2z-1)
where u – initial velocity, A -acceleration of the particle
 
a-b=A(x-y)
b-c=A(y-z)
(a-b)/(b-c)=(x-y)/(y-z)
ay-az-by+bz=bx-by-cx+cy
a(y-z)+b(z-x)+c(x-y)=0